OpenStudy (anonymous):

Complete the following: (a)           Use the Leading Coefficient Test to determine the graph's end behavior. (b)           Find the x-intercepts. State whether the graph crosses the x-axis or touches the x-axis and turns around at each intercept. Show your work. (c)           Find the y-intercept. Show your work. f(x) = x2(x + 2) (a).inf (b).touches at 0 and goes through -2 (c).2

OpenStudy (anonymous):

@phi my answers are up there but my teacher said this You need to provide the end behavior. On the right, is it going up or down and also what happens on the left? Distribute the x^2 before determining the degree of the polynomial. Then use the Leading Coefficient Test. Refer to section 2-6 for examples and sample problems, here you will find how to describe end behavior. 4b. At the x intercept, y=0. Since this is quadratic, how would you know if it intercepts or touches the x-axis?

OpenStudy (anonymous):

@Vocaloid

OpenStudy (loser66):

the cube function ( exponent 3) with positive leading coefficient has the left end at \(-\infty\) and right end at \(+ \infty\)

OpenStudy (loser66):

Hence, by looking at the leading coefficient ( the number stands right in the front of x^3) , you can see how the end looks like

OpenStudy (anonymous):

so how would i put it into words though

OpenStudy (loser66):

Put as what I wrote above.

OpenStudy (anonymous):

ohh ok

OpenStudy (loser66):

\(f(x) = x^2(x+2)=x^3+2x^2\) The leading coefficient is 1 >0 , hence the graph will start at -inf and end at + inf Dat sit

OpenStudy (anonymous):

1>0

OpenStudy (anonymous):

brb

OpenStudy (loser66):

to b) x intercepts happen when y =0, that is \(x^2 (x+2) =0\) , solve for x, you will have 2 solutions. Those are x-intercepts.

OpenStudy (loser66):

y-intercepts happen when x =0, replace x =0 to solve for y